# Uncertainty-enabled machine learning for emulation of regional sea-level change caused by the Antarctic Ice Sheet

### v1

@@ Abstract
Projecting sea-level change in various climate-change scenarios typically involves running forward simulations of the Earth’s gravitational, rotational and deformational (GRD) response to ice mass change, which requires high computational cost and time. Here we build neural-network emulators of sea-level change at 27 coastal locations, due to the GRD effects associated with future Antarctic Ice Sheet mass change over the 21st century. The emulators are based on datasets produced using a numerical solver for the static sea-level equation and published ISMIP6-2100 ice-sheet model simulations referenced in the IPCC AR6 report. We show that the neural-network emulators have an accuracy that is competitive with baseline machine learning emulators. In order to quantify uncertainty, we derive well-calibrated prediction intervals for simulated sea-level change via a linear regression postprocessing technique that uses (nonlinear) machine learning model outputs, a technique that has previously been applied to numerical climate models. We also demonstrate substantial gains in computational efficiency: a feedforward neural-network emulator exhibits on the order of 100 times speedup in comparison to the numerical sea-level equation solver that is used for training.
Keywords Regional sea level change  $\cdot$ Emulator  $\cdot$ Machine learning
@@ 1 Introduction
Sea-level change is a major impact of climate change that threatens coastal communities and infrastructure (Fox-Kemper et al.,, 2021; Hauer et al.,, 2020). Global mean sea level (GMSL) rose by 0.20 m over the last century, and the recent Intergovernmental Panel on Climate Change report (IPCC AR6; Fox-Kemper et al., (2021)) indicates that it is "virtually certain" that GMSL will continue to rise through 2100, regardless of anthropogenic greenhouse gas emissions scenario, due to the likelihood of all major processes contributing to GMSL continuing to operate. Sea-level change varies spatially, and relative sea level (RSL, i.e., ocean depth) can differ by up to a factor of two relative to GMSL across regions of the globe (Fox-Kemper et al.,, 2021; Kopp et al.,, 2015). Variations in RSL are caused by steric effects (changes in ocean density and salinity) and dynamic sea level changes, as well as gravitational, rotational, and deformational (GRD) effects (Slangen et al.,, 2014; Kopp et al.,, 2015; Gregory et al.,, 2019; Hamlington et al.,, 2020; Fox-Kemper et al.,, 2021). Unlike the other two factors, GRD effects are induced by changes in surface mass loading such as ice sheets, glaciers, and terrestrial water storage (reservoirs, lakes, groundwater). That is, mass redistribution on the Earth’s surface perturbs the gravitational field and rotation potential of the Earth. In addition, the solid Earth deforms over time in response to the changing surface loads. These complex processes yield spatially and temporally varying sea-level change referred to as "sea-level fingerprints" unique to the geometry of ice mass change (Farrell and Clark,, 1976; Mitrovica et al.,, 2011). Future sea-level change and uncertainties in projections will be dominated by changes in the Antarctic Ice Sheet (AIS) (Kopp et al.,, 2017; Bakker et al.,, 2017; Edwards et al.,, 2021; Fox-Kemper et al.,, 2021), making the sea-level change due to GRD effects an important factor to accurately predict.
Projections of future AIS contribution to sea level are improving through coordinated community model intercomparison efforts (Nowicki et al.,, 2020; Seroussi et al.,, 2020; Edwards et al.,, 2021), as well as large ensembles of AIS simulations from single models that quantify uncertainties (Bulthuis et al.,, 2019; Berdahl et al.,, 2023; Coulon et al.,, 2024). These ice-sheet projections have been used in probabilistic frameworks for analyzing regional sea level (Jevrejeva et al.,, 2018; Kopp et al.,, 2023). However, these frameworks have so far assumed static sea-level fingerprints calculated by only accounting for the total ice thickness changes for each ice sheet and glacier region (Kopp et al.,, 2023). One could improve the projections by considering temporally evolving sea-level fingerprint, as recent studies that predict sea-level changes due to GRD effects associated with projected land-ice mass change for 2015-2100 CE have shown that the variable geometry of AIS mass changes has a significant impact on future sea-level change (Roffman et al.,, 2023; Cederberg et al.,, 2023). Because probabilistic sea-level change frameworks require evaluation of tens of thousands of samples (Wong et al.,, 2017; Kopp et al.,, 2023), they typically rely on fast-to-evaluate emulators or parameterizations of complex models and processes. Tools for rapidly calculating sea-level fingerprints due to GRD effects for different patterns of land-ice mass change would facilitate probabilistic sea-level projections.
Because of the complex relationships between mass loading changes and sea-level fingerprints (Mitrovica et al.,, 2011; Roffman et al.,, 2023; Cederberg et al.,, 2023) and costs of calculating time-evolving sea-level fingerprints using sea-level models, machine learning (ML) provides a promising avenue for creating computationally inexpensive emulators of the sea-level equation. Recent work has demonstrated the potential of a graph-based spherical convolutional neural network to emulate a sea-level model applied to ice-age glacial isostatic adjustment over the last deglaciation and achieved a 100–1000$\times$ speedup (Lin et al.,, 2023). Another study used an artificial neural network to develop an emulator for the effect of lateral variations in Earth rheological structure that are often left out of sea-level models (Love et al.,, 2023). ML has also been applied to the ice-sheet model projections that form the inputs to sea-level models. Van Katwyk et al., (2023) created an emulator for the contribution from the AIS to GMSL by training a neural network on an ensemble of ice-sheet model experiments. Rohmer et al., (2022) used a similar ensemble, in this case for the Greenland Ice Sheet, in conjunction with an ML-based attribution approach to quantify the influence of different ice-sheet modeling decisions to projections of future ice-sheet contribution to GMSL. ML methods have also been utilized to predict complex regional sea level signals from observed historical sea-level change datasets at month to year timescales (e.g. Tur et al.,, 2021; Ayinde et al.,, 2023; Nieves et al.,, 2021; Guillou and Chapalain,, 2021; Altunkaynak and Kartal,, 2021; Song et al.,, 2021) and applied to various individual components of the ice-sheet models (e.g. Brinkerhoff et al.,, 2021; Jouvet et al.,, 2022; Jouvet,, 2023; He et al.,, 2023; Riel et al.,, 2021; Riel and Minchew,, 2023; Rosier et al.,, 2023; Verjans and Robel,, 2024). However, the use of ML to accelerate sea-level fingerprint calculations for projections of future regional sea level has yet to be attempted.
@@
Here, we develop a set of neural-network ML emulators (sometimes referred to as surrogate models) of the sea-level equation for predicting regional sea-level changes caused by changes in the mass of AIS and compare their performances to baseline Gaussian process and random forest based emulators. The emulators are trained on a set of sea-level model simulations forced with a recent ensemble of ice-sheet model projections of AIS through 2100 (ISMIP6, Serrousi et al., 2020). Rather than predicting the entire global field of sea-level change, we train the emulators to only consider the locations of 27 major coastal cities. In this paper, we begin by comparing four different ML methods for building emulators and evaluate their strengths and weaknesses for this application. We then develop an approach for quantifying uncertainty in the ML emulators using a simple linear regression technique to estimate prediction intervals of each emulator for each city, a method that has been used for calibrating numerical climate models (Glahn and Lowry,, 1972; Glahn et al.,, 2009; Lovegrove and Siegert,, 2023). Finally, we provide an illustration of applying the emulators to generate probabilistic projections of regional sea-level change caused by AIS mass changes by generating a large ensemble of artificial ice-sheet projections.
@@ 2 Data description
Training the ML emulators required datasets of ice-sheet mass change projections and the corresponding regional sea-level fingerprints, which were divided into training, validation and test groups. The following sections describe the generation of these datasets in detail.
@@ 2.1 Ice-sheet mass change dataset
The input dataset for our workflow was an ensemble of future AIS mass change calculated by ice-sheet models. For this, we adopted the results of the AIS experiments from the Ice Sheet Model Intercomparison Project for Coupled Model Intercomparison Project Phase 6 (ISMIP6) (Seroussi et al.,, 2020). The ISMIP6-Antarctica protocol specified standard experiment configurations for ice-sheet models to generate projections of AIS from 2015 through 2100 (Nowicki et al.,, 2020). The ice-sheet model simulations were forced by atmosphere and ocean output from six different Coupled Model Intercomparison Project Phase 5 (CMIP5) climate models for the high greenhouse gas emissions RCP8.5 scenario and two climate models for the low emissions RCP2.6 scenario. In addition to testing different climate forcings, different experiments tested different representations of ice-shelf basal melt and the potential for removal of ice-shelf regions through hydrofracture. In total, there were 21 experiments defined; here we select eight experiments for which standardized protocols were prescribed. The experiments were executed by 13 modeling groups around the world, and the results were analyzed on a common reference grid. Each modeling group also performed a control run through 2100 using steady constant climate forcing to quantify model drift. The ensemble represents a wide variety of ice-sheet model complexity, grid resolution, and model characteristics. Notably, none of the ice-sheet models in the ISMIP6 ensemble employed any of GRD effects.
@@
To prepare these ice-sheet projection data as input to a sea-level model, we performed a number of processing steps: masking out floating ice and removing model drift from projections on the 8 km-resolution standard polar stereographic grid, and interpolating the data onto a global Gaussian grid of spherical Earth. Since changes in floating ice do not contribute to sea level (except by a negligible amount due to density difference between ice and ocean water), ice thickness input to the sea-level model should only comprise grounded ice. We calculated changes in the ISMIP6 spatial field “land_ice_thickness” from its initial value. We restricted the calculation to the region of the domain where ice is grounded, which we calculate based on the flotation criterion, assuming an ice density of 910 kg m-3 and an ocean density of 1028 kg m-3. Second, because there is a large drift in the control run of many simulations (Aschwanden et al.,, 2021), we removed the thickness change in the control run for each model for each projection, as was done in the ISMIP6 analysis (Seroussi et al.,, 2020). This step is particularly important for modeling regional sea-level changes, as some portion of ice thickness change from raw projections will be related to the bias inherent to each ice-sheet model (i.e., model drift), which we do not want to feed into our sea-level fingerprint calculation. While ISMIP6 output is posted annually, we subsampled it to 5 year intervals because annual changes are relatively small, leading to 17 time levels of grounded ice thickness change. Finally, we remapped the ice thickness change fields from the 8 km polar stereographic projection to the $512\times 1024$ latitude/longitude Gaussian grid used by the sea-level model using a conservative remapping method to ensure total AIS mass change is unaltered by the remapping.
@@
A total of 194 AIS projections were generated for ISMIP6 across the different experiments and models used, and here we restrict ourselves to 71 projections derived from the eight experiments with standardized protocol (experiments 5-10 and 11-12), eliminating a small number of projections due to processing errors related to issues with data format conventions. Accounting for the 17 time levels retained for each projection, we had a total of 1207 fields of grounded ice thickness change in our input dataset used for training, validation, and testing of ML emulators.
@@ 2.2 Sea-level dataset: Calculation of sea-level fingerprints
We generated sea-level fingerprints associated with different patterns of AIS mass loss using the sea-level model described by Han et al., (2022), an updated version of sea-level models described in Kendall et al., (2005) and Gomez et al., (2010). The model computes barystatic sea-level changes due to GRD effects associated with AIS mass changes by solving the static sea-level equation (Farrell and Clark,, 1976; Mitrovica and Milne,, 2003) using the pseudo-spectral algorithm described in detail in Kendall et al., (2005), with truncation up to spherical harmonic degree and order 512. The sea-level framework accounts for shoreline migration, including inundation of presently glaciated marine basins as the extent of the grounded ice sheet retreats, as well as perturbations in the Earth’s rotation axis (Mitrovica and Milne,, 2003; Mitrovica et al.,, 2005).
@@
In addition to the projected patterns of AIS ice mass loss, the simulation requires a model for the rheological structure of the solid Earth. We adopt a radially symmetric Maxwell viscoelastic model for the Earth with elastic and density properties derived from the seismic PREM model (Dziewonski and Anderson,, 1981). The depth varying viscosity of the Earth is given by the viscosity profile VM7, which is based on recent inferences from wide range of global geophysical observables (Roy and Peltier,, 2017, 2018). We note that these values represent a global average; while appropriate for East Antarctica and our 27 far field coastal locations, this Earth structure does not accurately capture the Earth structure in West Antarctica where upper mantle viscosity is thought to be 2-3 orders of magnitude lower than the global average (Barletta et al.,, 2018; Nield et al.,, 2014). We also note that while it is well-established that the Earth structure in Antarctica shows strong lateral heterogeneity (Lloyd et al.,, 2023) likely having an effect on projections of local sea-level change, the sensitivity of far-field sea level to the 1D versus 3D Earth structure only becomes large enough to consider in a long-term (multi-centennial or longer) scale projection (Pan et al.,, 2021; Powell et al.,, 2022), which is beyond the timescale we consider in this work. While we include viscoelastic deformation of the solid Earth in the model, with the chosen viscosity profile over centennial timescales, the viscous component is significantly smaller than the elastic component, yielding results similar to assuming an elastic Earth.
@@
Since the sea-level model runs on a global grid but the original ISMIP6-AIS datasets were provided on a local grid, we adopt topography from the ETOPO2 dataset (NOAA National Centers for Environmental Information,, 2022) for the region outside of the ice-sheet model domain. With the pre-processed ice-sheet thickness change time-series and the Earth structure described above, we perform a suite of sea-level simulations with 5-year time intervals and generate sea-level change output datasets for emulator training, validation, and testing. Examples of the maps of sea-level change associated with AIS thickness changes are illustrated in Figure 1.
@@
We initially attempted to emulate the entire $512\times 1,024=524,288$ dimension output space of the sea-level model, with the best-performing model using an autoencoder (Kramer,, 1991) for dimension reduction on inputs and outputs and a feed-forward neural network linking between reduced-dimension spaces. Our initial results varied considerably in quality – while some of the entire fields in the test set could be emulated reliably, about 10% of the global test fields were poorly emulated (normalized root mean squared error exceeding 0.11). Moreover, the latitude-longitude coordinates of the sea-level model spatial output added complexity in accurately representing a spherical surface in two dimensions (Lin et al.,, 2023). To create a more tractable but useful demonstration case, we chose to only evaluate RSL change at a subset of coastal locations. We chose the 27 long-term tide gauge sites identified by Meyssignac et al., (2017) (Fig. 1c), which come from the Permanent Service for Mean Sea Level (Holgate et al.,, 2013).
@@
Figure 1: Illustration of RSL fingerprints for given AIS thickness changes for two samples. a) Modeled AIS thickness change between 2015 and 2100 for sample from training dataset with third largest total AIS mass change. b) Calculated RSL change fingerprint for AIS mass loss in a. c) Modeled AIS thickness change between 2015 and 2100 for sample from training dataset with third smallest total AIS mass change. d) Calculated RSL change fingerprint for AIS mass loss in c. Note different colorbar ranges for each plot and that values exist beyond the ranges shown. The 27 long-term tide gauge sites selected by Meyssignac et al., (2017) that we used as emulation targets are shown with black dots in b and d.

### v2

# Emulation with uncertainty quantification of regional sea-level change caused by the Antarctic Ice Sheet

@@ Abstract
Projecting regional sea-level change under various climate-change scenarios typically involves running forward simulations of the Earth’s gravitational, rotational and deformational (GRD) response to ice-mass change, which requires substantial computational cost if applied to probabilistic frameworks requiring thousands to millions of samples. Here we build emulators of regional sea-level change at 27 coastal locations, due to the GRD effects associated with future Antarctic Ice Sheet mass change over the 21st century. The emulators are evaluated against a numerical sea-level model applied to an ensemble of ice-sheet model simulations of the Antarctic Ice Sheet through 2100. We build a physics-based emulator using a recent sensitivity kernel approach and compare it to machine learning based emulators (neural network and conditional variational autoencoder methods). In order to quantify uncertainty, we derive well-calibrated prediction intervals for regional sea-level change via split-conformal inference and linear regression, and show that Monte Carlo dropout does not yield well-calibrated uncertainties in this instance. We also demonstrate substantial gains in computational efficiency using both the physics-based emulator and neural networks in comparison to the numerical model for the complete regional sea-level solution. Overall, we find the physics-based emulator modestly outperforms the machine learning emulators for this problem.
Plain Language Summary
Sea-level change varies regionally because changes in the weight of ice sheets and glaciers deform the Earth’s crust and modify the Earth’s gravitational field and axis of rotation. While numerical sea-level models are widely used to project sea-level change from these effects, they require thousands of evaluations to quantify uncertainty in the projections. To accelerate this process, we evaluate the feasibility of different simplified models that trade some accuracy for substantial computational speedup. We consider both machine learning methods and a reduced-physics approach and apply these “emulators” to 27 major coastal locations. The emulators demonstrate high prediction accuracy and are between 160 - 240 times faster to run than the numerical sea-level model. We then quantify the uncertainty in regional sea-level change introduced by the emulators themselves using two techniques. Based on this, we find that the reduced-physics approach is most accurate, fast to evaluate, is easy to generate, and produces accurate uncertainties when coupled with the uncertainty quantification techniques aforementioned.
@@ 1 Introduction
Coastal communities and infrastructure are under severe threat from sea-level rise caused by climate change (foxkemper2021ocean; Hauer2020a). Global mean sea level (GMSL) rose by 0.20 m over the last century, and the most recent Intergovernmental Panel on Climate Change assessment report (foxkemper2021ocean) indicates that it is "virtually certain" that GMSL will continue to rise through 2100. Continued sea-level rise is expected regardless of anthropogenic greenhouse gas emissions scenario due to the likelihood of all major processes contributing to GMSL continuing to operate. Sea-level change varies spatially, and relative sea level (RSL, i.e., thickness of the ocean) can differ by up to a factor of two relative to GMSL across regions of the globe (foxkemper2021ocean; Kopp2015). Variations in RSL are caused by steric effects (changes in ocean density and salinity) and dynamic sea level changes, as well as gravitational, rotational, and deformational (GRD) effects (Slangen2014; Kopp2015; Gregory2019; Hamlington2020; foxkemper2021ocean). Unlike the other two factors, GRD effects are induced by changes in surface mass loading such as ice sheets and ocean, glaciers, and terrestrial water storage (reservoirs, lakes, groundwater). That is, mass redistribution on the Earth’s surface deforms the solid Earth, perturbs the Earth’s gravitational field, and shifts the rotation axis of the Earth. In addition, the solid Earth deforms over time in response to the changing surface loads. These complex processes yield spatially and temporally varying sea-level changes referred to as "sea-level fingerprints" unique to the geometry of ice mass change (Farrell1976; Mitrovica2011).
Uncertainty in future sea-level change is dominated by changes in the Antarctic Ice Sheet (AIS) (Kopp2017; Bakker2017; Edwards2021; foxkemper2021ocean), because it contains large sectors where ice is grounded below sea level and potentially susceptible to rapid retreat through marine ice-sheet instability (Schoof2007). Projections of future AIS contribution to sea level are improving through coordinated community model intercomparison efforts (Nowicki2020; Seroussi2020; Edwards2021), as well as large ensembles of AIS simulations from single models that quantify uncertainties (Bulthuis2019; Berdahl2023; Coulon2024). These studies indicate a wide range of potential future outcomes for AIS through 2100 and beyond. Further, recent studies have shown that the variable geometry of projected AIS mass changes has a significant impact on future RSL due to GRD effects (Roffman2023; Cederberg2023).
Because sea-level change has large uncertainty and also varies regionally, probabilistic projections of RSL at cities of interest that incorporate the full range of AIS projections are critical for assessing future climate impacts and risk. These ice-sheet projection ensembles have been used in probabilistic frameworks for analyzing RSL that use Monte Carlo sampling with large numbers of evaluations (order $10^{5}-10^{6}$) to produce probability distribution functions of regional sea level at individual cities of interest (Jevrejeva2018; Kopp2023). However, these frameworks have so far assumed static sea-level fingerprints calculated by only accounting for the total ice thickness changes for each ice sheet and glacier region (Kopp2023). These projections would be improved by considering temporally evolving sea-level fingerprints. Because probabilistic sea-level change frameworks require evaluation of tens of thousands of samples (Wong2017; Kopp2023), they typically rely on fast-to-evaluate emulators or parameterizations of complex models and processes that include an acceptable trade-off between accuracy and computational cost. Our goal in this work is to evaluate two methods for introducing sample-specific sea-level fingerprints into the probabilistic frameworks that produce RSL distributions for individual cities: emulation of a sea-level model using 1. machine learning, and 2. sensitivity kernels (Al-Attar2024) derived from an adjoint method for a simplified version of the sea-level model.
An emulator is a surrogate model that mimics the underlying simulator (such as the sea-level model in this study) and is computationally efficient, serving as an alternative to the model it replicates. The significance of an emulator becomes particularly evident when the simulator is computationally too expensive and when the emulator can effectively quantify uncertainties. Because of the complex relationships between mass loading changes and sea-level fingerprints (Mitrovica2011; Roffman2023; Cederberg2023) and costs of calculating time-evolving sea-level fingerprints using sea-level models, machine learning (ML) provides one promising avenue for creating computationally inexpensive emulators of sea-level fingerprints evaluated at locations of interest.
[…]
The second approach we evaluate employs a sensitivity kernel for each location of interest (city) against which fields of ice mass change can be multiplied to derive RSL for that location without needing to re-solve the sea-level equation (Al-Attar2024). The sensitivity kernel approach exploits recently established reciprocity theorems for the elastic sea-level equation that show that the RSL at one location due to the change in mass load (e.g., from an ice sheet) at another location is equal if the loading and RSL evaluation points are reversed (Al-Attar2024). The kernel approach requires some minor limiting assumptions (elastic rheology; fixed shoreline, including ice-sheet grounding lines; inability to self-consistently evaluate the ice-sheet flotation condition) be applied to the sea-level equation and must be calculated separately for each location of interest. These restrictions motivate the comparison with ML emulation as a potential alternative.
@@
In this study, we compare neural-network ML emulation approaches against the kernel sensitivity approach for the purpose of rapid evaluation of many samples at a modest number of geographic locations. We develop multiple neural-network ML emulators for the established sea-level equation solver of Han2022 for predicting RSL changes caused by changes in the mass of AIS. The emulators are trained on a set of sea-level model simulations forced with a recent ensemble of ice-sheet model projections of AIS through 2100 (ISMIP6, Serrousi et al., 2020). Because we are focused on methods for producing RSL distributions for individual cities, we train the emulators to consider the locations of 27 major coastal cities; note that the neural-network emulators jointly predict RSL changes for all 27 cities at once, whereas the kernel sensitivity approach has a distinct kernel for each of the 27 locations. We begin by comparing different neural-network based emulators, as well as the kernel sensitivity method, and evaluate their strengths and weaknesses for this application. We then quantify uncertainty in emulator predictions by generating prediction intervals using both split-conformal inference (lei2018distribution) and simple linear regression, a method that has been used for calibrating numerical forecasting models (glahn1972use; glahn2009mos; lovegrove2023improving). We also compare Monte Carlo dropout (dropout), a commonly-used method for uncertainty quantification in neural networks. Next, we apply the emulators to generate probabilistic projections of RSL change caused by AIS mass changes by generating an ensemble of artificial ice-sheet projections. Finally, we assess the relative cost and accuracy of the ML emulation and kernel sensitivity approaches for this application.
@@ 2 Data description
Training the ML emulators required datasets of simulated ice-sheet mass change projections and the corresponding regional sea-level fingerprints, which were divided into training, validation and test groups. The following sections describe the generation of these datasets in detail.
@@ 2.1 Ice-sheet mass change dataset
The input dataset for our workflow was an ensemble of future AIS mass change calculated (i.e., simulated) by ice-sheet models. For this, we adopted the results of the AIS experiments from the Ice Sheet Model Intercomparison Project for Coupled Model Intercomparison Project Phase 6 (ISMIP6) (Seroussi2020). The ISMIP6-Antarctica protocol specified standard experiment configurations for ice-sheet models to generate projections of AIS from 2015 through 2100 (Nowicki2020). The ice-sheet model simulations were forced by atmosphere and ocean output from six different Coupled Model Intercomparison Project Phase 5 (CMIP5) climate models for the high greenhouse gas emissions RCP8.5 scenario and two climate models for the low emissions RCP2.6 scenario. In addition to testing different climate forcings, different experiments tested different representations of ice-shelf basal melt and the potential for removal of ice-shelf regions through hydrofracture. In total, there were 21 experiments defined; here we select eight experiments for which standardized protocols were prescribed. The experiments were executed by 13 modeling groups around the world, and the results were analyzed on a common reference grid. Each modeling group also performed a control run through 2100 using steady constant climate forcing to quantify model drift. The ensemble represents a wide variety of ice-sheet model complexity, grid resolution, and model characteristics. Notably, none of the ice-sheet models in the ISMIP6 ensemble included GRD effects.
@@
To prepare these ice-sheet projection data as input to a sea-level model, we performed a number of processing steps. First, we calculated changes in the ISMIP6 spatial field “land_ice_thickness” from its initial value. Because there is a large drift in the control run of many simulations (Aschwanden2021), we removed the thickness change in the control run for each model from each projection, as was done in the ISMIP6 analysis (Seroussi2020). This step is particularly important for modeling regional sea-level changes, as some portion of ice thickness change from raw projections will be related to the bias inherent to each ice-sheet model (i.e., model drift), which we do not want to feed into our sea-level fingerprint calculation. While ISMIP6 output is posted annually, we subsampled it to 5 year intervals because annual changes are relatively small, leading to 17 time levels of ice thickness change. Finally, we remapped the ice thickness change fields from the 8 km polar stereographic projection to the $512\times 1024$ latitude/longitude Gaussian grid used by the sea-level model using a conservative remapping method to ensure total AIS mass change is unaltered by the remapping.
@@
A total of 194 AIS projections were generated for ISMIP6 across the different experiments and models used, and here we restrict ourselves to 71 projections derived from the eight experiments with standardized protocol (experiments 5-10 and 11-12), eliminating a small number of projections due to processing errors related to issues with data format conventions. Accounting for the 17 time levels retained for each projection, we had a total of 1207 fields of ice thickness change in our input dataset used for training, validation, and testing of ML emulators.
@@ 2.2 Sea-level dataset: Calculation of sea-level fingerprints
We generated sea-level fingerprints associated with different patterns of AIS mass loss using the sea-level model described by Han2022, an updated version of sea-level models described in Kendall2005 and Gomez2010a. The model computes barystatic sea-level changes and associated GRD effects caused by land ice mass changes by solving the static sea-level equation (Farrell1976; Mitrovica2003) using the pseudo-spectral algorithm described in detail in Kendall2005, with truncation up to spherical harmonic degree and order 512. The sea-level framework accounts for shoreline migration (Milne_Mitrovica_1998), including inundation of presently glaciated marine basins as the extent of the grounded ice sheet retreats, as well as perturbations in the Earth’s rotation axis (Mitrovica2003; Mitrovica2005). Our simulations include a check for floating ice by evaluating the flotation criterion iteratively as RSL is solved. The flotation criterion assumes an ice density of 910 kg m-3 and a water (ice melt and ocean) density of 1000 kg m-3. Note that because we are interested in projections of future sea-level change, there are not observational data to evaluate our sea-level change calculations against and we assume that the calculations by the Han2022 model are sufficiently accurate.
@@
In addition to the projected patterns of AIS ice mass loss, the simulation requires a model for the rheological structure of the solid Earth. We adopt a radially symmetric elastic model for the Earth with elastic and density properties derived from the seismic PREM model (Dziewonski1981). We also evaluated simulations with viscoelastic Earth structure using the depth-varying viscosity of the Earth given by the viscosity profile VM7 (Roy2017; Roy2018). For the 85 year time-frame of the ISMIP6 experiments the contribution from viscous deformation was negligible. We note that these viscosity values represent a global average; while appropriate for East Antarctica and our 27 far field coastal locations, this Earth structure does not accurately capture the Earth structure in West Antarctica where upper mantle viscosity is thought to be 2-3 orders of magnitude lower than the global average (Barletta2018; Nield2014). We note that while it is well-established that the Earth structure in Antarctica shows strong lateral heterogeneity (Lloyd2020) likely having an effect on projections of local sea-level change, the sensitivity of far-field sea level to the 1D versus 3D Earth structure only becomes large enough to consider in a long-term (multi-centennial or longer) scale projection (Pan2021; Powell2022), which is beyond the timescale we consider in this work. We conducted additional sensitivity tests using a lower mantle viscosity appropriate for West Antarctica and confirmed that results differed only by a few percent from the elastic-Earth configuration outside the Antarctic for these 85-year simulations, justifying our use of an elastic-only Earth structure.
@@
Since the sea-level model runs on a global grid but the original ISMIP6-AIS datasets were provided on a local grid, we adopt topography from the ETOPO2 dataset (NOAA2022) for the region outside of the ice-sheet model domain. With the pre-processed ice-sheet thickness change time-series and the Earth structure described above, we perform a suite of sea-level simulations with 5-year time intervals and generate sea-level change output datasets for emulator training, validation, and testing. Examples of the maps of sea-level change associated with AIS thickness changes are illustrated in Figure 1. Note that while the 17 time levels evaluated are sequential in time, the elastic GRD effects considered are instantaneous and there is no time-dependence in the model solutions. That is, the RSL field is calculated from a given ice thickness change field with no dependence on ice thickness change history.
@@
We initially attempted to emulate the entire $512\times 1,024=524,288$ dimension output space of the sea-level model (per time step), with the best-performing model using an autoencoder (autoencoder) for dimension reduction on inputs and outputs and a feed-forward neural network linking between reduced-dimension spaces. Our initial results varied considerably in quality – while some of the entire fields in the test set could be emulated reliably, about 10% of the global test fields were poorly emulated (normalized root mean squared error exceeding 0.11). Moreover, the latitude-longitude coordinates of the sea-level model spatial output added complexity in accurately representing a spherical surface in two dimensions (lin2023georgia). Given our primary goal of emulating RSL at a set of selected cities, the global RSL field was not strictly necessary. Therefore, to improve the accuracy of the ML emulators, we chose to only evaluate RSL change at the 27 long-term tide gauge sites identified by Meyssignac2017a (Fig. 1c), which come from the Permanent Service for Mean Sea Level (Holgate2013).
@@
Figure 1: Illustration of RSL fingerprints for given AIS thickness changes for two samples representative of small and large AIS mass change. a) Modeled AIS grounded-ice thickness change between 2015 and 2100 for sample from training dataset with third smallest total AIS mass change. b) Calculated RSL change fingerprint for AIS mass loss in a. c) Modeled AIS grounded-ice thickness change between 2015 and 2100 for sample from training dataset with third largest total AIS mass change. d) Calculated RSL change fingerprint for AIS mass loss in c. Note different colorbar ranges for each plot and that values exist beyond the ranges shown. The 27 long-term tide gauge sites selected by Meyssignac2017a that we used as emulation targets are shown with black dots in b and d.